# e-Profits: A Business-Aligned Evaluation Metric for Profit-Sensitive Customer Churn Prediction

### v1

@@ Abstract
Retention campaigns in customer relationship management often rely on churn prediction models evaluated using traditional metrics such as AUC and F1-score. However, these metrics fail to reflect financial outcomes and may mislead strategic decisions. We introduce e-Profits, a novel business-aligned evaluation metric that quantifies model performance based on customer-specific value, retention probability, and intervention costs. Unlike existing profit-based metrics such as Expected Maximum Profit, which assume fixed population-level parameters, e-Profits uses Kaplan-Meier survival analysis to estimate personalised retention rates and supports granular, per-customer evaluation. We benchmark six classifiers across two telecom datasets (IBM Telco and Maven Telecom) and demonstrate that e-Profits reshapes model rankings compared to traditional metrics, revealing financial advantages in models previously overlooked by AUC or F1-score. The metric also enables segment-level insight into which models maximise return on investment for high-value customers. e-Profits is designed as an understandable, post hoc tool to support model evaluation in business contexts, particularly for marketing and analytics teams prioritising profit-driven decisions. All source code is available at: https://github.com/matifq/eprofits.
@@
Churn prediction , Machine learning & Artificial Intelligence , Profitable maximizing churn prediction , Profit based metrics
journal: Journal of Retailing and Consumer Services
[…]
@@ 1 Introduction
Customer churn, the end of a relationship between a customer and a business, presents a serious risk to a company’s sustainability. Whether customers choose to go to a competitor or stop using the service altogether, churn results in immediate revenue loss and a gradual decline in customer equity. Since attracting new customers is much more expensive typically five times more than keeping current ones [1, 2], many companies now focus on customer retention strategies enhanced by predictive analytics [3, 4]. Loyal customers may gradually acquire more goods and services over time and are associated with reduced marketing costs [5], and they spread positive sentiment which indirectly attract new customers. Therefore, higher customer retention rates result in increases in company profits .
@@
Businesses increasingly rely on predictive models trained on behavioural and transactional data to address churn and identify customers at risk of leaving [6, 7]. Customer churn prediction (CPP) is typically framed as a binary classification task in the machine learning literature [8]. The goal is to identify would-be churners early enough to intervene, usually via targeted retention campaigns offering incentives. However, while predictive accuracy is important, most models are evaluated using traditional metrics such as F1-score or AUC, which do not reflect real-world business impact [9, 10]. This issue is especially problematic in profit-sensitive domains, where not all customers are equally valuable, and not all classification errors have the same cost [11, 12]. For instance, targeting low-value customers who were unlikely to churn wastes retention resources, while failing to identify high-value churners results in substantial lost revenue. Studies have shown that an inverse relationship may exist between customer lifetime value (CLV) and churn probability [13, 14], meaning that models can systematically overlook high-value customers optimised for statistical accuracy.
@@
To bridge this gap, profit-based evaluation metrics have been proposed, including the Maximum Profit Criterion (MPC) and Expected Maximum Profit (EMP) [15, 16]. These metrics represent important progress but are often constrained by fixed assumptions, such as uniform CLV and constant offer acceptance probabilities, which rarely hold in real-world business contexts characterised by customer heterogeneity.
@@
In this study, we introduce e-Profits, a business-aligned evaluation metric designed to assess churn models based on financial relevance rather than statistical precision. By incorporating customer-specific CLV, intervention cost, and retention probability, estimated through Kaplan-Meier survival analysis, e-Profits offers a granular, per-customer evaluation approach. We empirically evaluate six machine learning classifiers across two real-world telecom datasets (IBM Telco and Maven Telecom) and show that e-Profits reveals model rankings that diverge significantly from those based on traditional or existing profit-based metrics.
@@ 2 Related Works
Customer churn prediction has been extensively studied across domains such as telecommunications, banking, and subscription-based services [17]. Traditionally, churn is formulated as a supervised binary classification problem, with models trained to distinguish between customers who stay and those who leave. Early approaches employed logistic regression and decision trees, gradually evolving toward advanced techniques, including artificial neural networks [18], support vector machines [19], ensemble methods [20], and feature selection pipelines [21]. Despite these advances, the dominant focus remains on improving predictive performance measured through conventional metrics such as Accuracy, F1-score, and AUC [22].
@@
These conventional metrics, however, are ill-suited for business settings where the cost of misclassification varies across customers. For example, AUC treats false positives and false negatives symmetrically [9], ignoring the financial impact of retention decisions. As Elkan noted in his foundational work on cost-sensitive learning [10], a classifier’s utility must be assessed based on domain-specific cost matrices rather than classification error alone. In churn prediction, misclassifying a high-value customer as a non-churner could lead to significant revenue loss, while misclassifying a low-value, loyal customer could result in wasted marketing resources.
@@
Profit-based evaluation metrics have been introduced to address this gap. Verbeke et al. introduced the Maximum Profit Criterion (MPC) [15], which estimates profit based on classification outcomes. This was extended by Verbraken et al. through the Expected Maximum Profit (EMP) metric [16], which incorporates uncertainty in retention success and introduces the profit-maximising churn fraction, the optimal subset of customers to target. These metrics marked an important shift from accuracy-focused to business-aligned evaluation. Table 1 summarises the key characteristics of traditional and profit-aware evaluation metrics, highlighting the business relevance and flexibility offered by e-Profits compared to widely used alternatives.
@@
Nevertheless, limitations persist. EMP assumes constant values for customer lifetime value (CLV), contact cost, and offer acceptance probability [23]. It models retention success using an unimodal Beta distribution with fixed hyperparameters, which may not capture the heterogeneity of real-world customers. Additionally, while some studies have explored embedding profit functions directly into model optimisation via profit-aware decision trees or evolutionary strategies [22, 24], these approaches often introduce complexity and reduce interpretability, limiting their uptake in business environments.
@@
Some recent works have considered time-dependent churn risk using survival analysis models [25, 26]. These approaches explicitly demonstrate the value of incorporating tenure and retention time, but they seldom translate into actionable profit-based evaluation metrics.
@@
Metric Type Business-Aware Fixed-Assumptions Individual-Level Optimisation Target
AUC Ranking No No Churn likelihood
F1-score Classification No No Class balance
EMP Profit-based Partial Yes No Profit-max subset
e-Profits Profit-based Yes No Yes Total and segment profit
While the academic literature contains several profit-aware metrics, their adoption in operational churn platforms remains rare. Surveys such as García et al. [27] and our own review [28] highlight that many practitioners still default to traditional metrics like AUC or lift, due to limitations in available tools and ease of understanding. This disconnect reinforces the need for practical and business-aligned evaluation metrics that can bridge the research-practice divide.
Conventional evaluation metrics such as Accuracy, F1-score, and AUC focus on statistical correctness but neglect the financial implications of churn prediction. Although profit-based metrics such as the EMP [16] represent an important step toward business-aligned evaluation, they often rely on population-level assumptions: fixed CLV, fixed intervention cost parameters (such as cost of contact and retention offer), and no integration of individualised retention rates. Such assumptions ignore the heterogeneity of real-world customer bases, where profit potential and intervention costs vary significantly across individuals.
Despite these advances, existing profit-aware approaches generally fail to capture customer heterogeneity. Most rely on fixed cost parameters or make strong assumptions about offer acceptance rates. By contrast, our proposed e-Profits is aimed at leveraging individual-level cost, revenue, and survival-based retention to provide a metric that is granular, understandable, and practically deployable, without retraining the model. e-Profits responds to these limitations by offering a modular evaluation metric that aligns directly with business impact, empowering practitioners to make financially sound model selection decisions.
@@ 3.1 Overview and Motivation
As established in the previous section, conventional evaluation metrics fail to capture the financial implications of churn misclassification, and existing profit-based alternatives often rely on strong population-level assumptions. In response, we introduce e-Profits, a modular evaluation metric that supports granular, business-aligned model comparison.
The e-Profits metric evaluates model performance by aggregating per-customer profit outcomes under realistic operational conditions. It incorporates:
  • Customer-specific lifetime value (CLV): Estimated using revenue and retention projections
  • Personalised retention probabilities: Derived via Kaplan-Meier survival analysis based on customer tenure
  • Individualised intervention costs: Including both contact and retention offer costs per customer
This design enables profit-based comparison of churn models without altering training pipelines, ensuring both understandability and practical deployment feasibility.
@@
Figure 1: Conceptual workflow of the e-Profits evaluation metric. This flow illustrates how model predictions interact with customer-level retention, CLV, and cost information to compute aggregate business impact. (To calculate CLV, we have used average and tenure-based retention rates (Using Kaplan-Meier))
@@
We define e-Profits as the total estimated profit that would be realised if a retention campaign was executed based on the model’s churn predictions:
@@
Where $\Pi_{i}$ is the estimated profit or loss associated with customer $i$, computed as:
$\Pi_{i}=\begin{cases}\text{CLV}_{i}-C_{\text{offer},i}-C_{\text{contact},i}&\text{if }y_{i}=1\text{ and }\hat{y}_{i}=1\quad\text{(True Positive)}\\ \\ -C_{\text{offer},i}-C_{\text{contact},i}&\text{if }y_{i}=0\text{ and }\hat{y}_{i}=1\quad\text{(False Positive)}\\ \\ 0&\text{otherwise (no retention action)}\end{cases}$ (2)
Where:
  • $y_{i}$ is the true churn label for customer $i$
  • $\hat{y}_{i}$ is the predicted label (1 = predicted to churn)
  • $\text{CLV}_{i}$ is the estimated customer lifetime value
  • $C_{\text{offer},i}$ and $C_{\text{contact},i}$ are the intervention costs for customer $i$
@@
3.3 Customer Lifetime Value (CLV)
The CLV for each customer is calculated using the revenue, the estimated retention probability, and a fixed profit margin:
$\text{CLV}_{i}=\frac{R_{i}\times M}{1-\min(r_{i},0.995)}$ (3)
Where:
  • $R_{i}$ is the monthly revenue from customer $i$
  • $M$ is the profit margin
  • $r_{i}$ is the customer-specific retention probability
The minimum function ensures numerical stability by capping the retention probability at 0.995. This avoids division by values close to zero, which could lead to unrealistically large CLV estimates, especially in cases where long-tenured customers receive extremely high retention probabilities from survival models like Kaplan-Meier.
3.4 Retention Rate Estimation via Survival Analysis
To estimate the probability that a customer remains subscribed beyond a given tenure, we employ Kaplan-Meier survival analysis [25]. This method is non-parametric, meaning it does not assume any specific distribution of churn times, making it well-suited for business datasets where churn dynamics are irregular or unknown [29]. Compared to parametric or neural survival models, Kaplan-Meier offers a transparent and computationally efficient approach [30] that facilitates easy integration into evaluation workflows such as e-Profits.

### v2

@@ Abstract
Retention campaigns in customer relationship management often rely on churn prediction models evaluated using traditional metrics such as AUC and F1-score. However, these metrics fail to reflect financial outcomes and may mislead strategic decisions. We introduce e-Profits, a novel business-aligned evaluation metric that quantifies model performance based on customer lifetime value, retention probability, and intervention costs. Unlike existing profit-based metrics such as Expected Maximum Profit, which assume fixed population-level parameters, e-Profits uses Kaplan-Meier survival analysis to estimate tenure-conditioned (customer-level) one-period retention probabilities and supports granular, per-customer profit evaluation. We benchmark six classifiers across two telecom datasets (IBM Telco and Maven Telecom) and demonstrate that e-Profits reshapes model rankings compared to traditional metrics, revealing financial advantages in models previously overlooked by AUC or F1-score. The metric also enables segment-level insight into which models maximise return on investment for high-value customers. e-Profits provides a transparent, customer-level evaluation framework that bridges predictive modelling and profit-driven decision-making in operational churn management. All source code is available at: https://github.com/Awaismanzoor/eprofits.
@@
Churn prediction, Machine learning & Artificial Intelligence, Profit maximizing churn prediction, Customer Relationship Management
@@ 1 Introduction
Customer churn, the end of a relationship between a customer and a business, presents a serious risk to a company’s sustainability. Whether customers choose to go to a competitor or stop using the service altogether, churn results in immediate revenue loss and a gradual decline in customer equity. Since attracting new customers is often reported to be more expensive than keeping current ones, with many studies and managerial sources suggesting that acquisition costs substantially exceed retention costs [van2004customer, zhou2023early], many companies focus on customer retention strategies enhanced by predictive analytics [ganesh2000understanding, sun2021case]. Loyal customers may gradually acquire more goods and services over time and are associated with reduced marketing costs [loureiro2025improving], and they spread positive sentiment, which indirectly attract new customers. Therefore, higher retention directly increases company profits.
@@
Businesses increasingly rely on predictive models trained on behavioural and transactional data to address churn and identify customers at risk of leaving [kotan2025novel, peng2023research]. Customer churn prediction (CPP) is typically framed as a binary classification task in the machine learning literature [faraji2024customer]. The goal is to identify would-be churners early enough to intervene, usually via targeted retention campaigns offering incentives. However, while predictive accuracy is important, most models are evaluated using traditional metrics such as F1-score or AUC, which do not reflect real-world business impact [hand2009measuring, elkan2001foundations]. This issue is especially problematic in profit-sensitive domains, where not all customers are equally valuable, and not all classification errors have the same cost [bahnsen2015novel, thakkar2022clairvoyant]. For instance, targeting low-value customers who were unlikely to churn wastes retention resources, while failing to identify high-value churners results in substantial lost revenue. Studies have shown that an inverse relationship may exist between customer lifetime value (CLV) and churn probability [rahman2025profit, lemmens2020managing], meaning that models can systematically overlook high-value customers optimised for statistical accuracy.
@@
To bridge this gap, profit-based evaluation metrics have been proposed, including the Maximum Profit Criterion (MPC) and Expected Maximum Profit (EMP) [verbeke2012new, verbraken2012novel]. These metrics represent important progress but are often constrained by fixed assumptions, such as uniform CLV and constant offer acceptance probabilities, which rarely hold in real-world business contexts characterised by customer heterogeneity.
@@
In this study, we introduce e-Profits, a business-aligned evaluation metric designed to assess churn models based on financial relevance rather than statistical precision. By incorporating customer-specific lifetime value (CLV), intervention costs, and retention probabilities estimated via Kaplan-Meier survival analysis, e-Profits enables granular, per-customer profitability assessment. We empirically evaluate six machine learning classifiers across two real-world telecom datasets (IBM Telco and Maven Telecom) and show that e-Profits reveals model rankings that diverge significantly from those obtained using traditional predictive metrics and existing profit-based alternatives. Beyond technical robustness, e-Profits aligns model evaluation with organisational decision-making by translating predictive outputs into financially actionable insights that are interpretable by both technical and non-technical stakeholders.
This paper introduces e-Profits, a business-aligned evaluation metric for customer churn prediction that enables model selection based on financial impact rather than predictive accuracy alone. Unlike conventional metrics such as AUC, Accuracy, and F1-score, and unlike existing profit-based metrics that rely on restrictive population-level assumptions, e-Profits provides a transparent, customer-level profit evaluation framework that can be applied to any trained classifier without modifying the learning algorithm.
The study makes the following key contributions:
  1. 1.
    Introduce e-Profits, a business-aligned evaluation metric for profit-sensitive churn prediction that integrates customer-level value, intervention costs, and retention probabilities;
  2. 2.
    Incorporate survival-based retention estimation using both average (ARR) and tenure-conditioned (TRR) Kaplan-Meier formulations;
  3. 3.
    Demonstrate that profit-aware evaluation induces statistically significant model re-ranking relative to traditional objectives such as AUC and F1-score;
  4. 4.
    Analyses both full-population and budget-constrained (top-segment) targeting scenarios relevant to real-world retention campaigns; and
  5. 5.
    Validates the robustness of profit-based model selection using bootstrap confidence intervals and paired non-parametric tests across two real-world datasets.
@@ 2 Related Works
Customer churn prediction has been extensively studied across domains such as telecommunications, banking, and subscription-based services [bhattacharyya2021investigation]. Traditionally, churn is formulated as a supervised binary classification problem, with models trained to distinguish between customers who stay and those who leave. Early approaches employed logistic regression and decision trees, gradually evolving toward advanced techniques, including artificial neural networks [mena2024exploiting], support vector machines [ly2022churn], ensemble methods [tavassoli2022hybrid], and feature selection pipelines [sana2022novel]. Despite these advances, the dominant focus remains on improving predictive performance measured through conventional metrics such as Accuracy, F1-score, and AUC [stripling2018profit].
@@
These conventional metrics, however, are ill-suited for business settings where the cost of misclassification varies across customers. For example, AUC treats false positives and false negatives symmetrically [hand2009measuring], ignoring the financial impact of retention decisions. As Elkan noted in his foundational work on cost-sensitive learning [elkan2001foundations], a classifier’s utility must be assessed based on domain-specific cost matrices rather than classification error alone. In churn prediction, misclassifying a high-value customer as a non-churner could lead to significant revenue loss, while misclassifying a low-value, loyal customer could result in wasted marketing resources.
@@
Profit-based evaluation metrics have been introduced to address this gap. Verbeke et al. introduced the Maximum Profit Criterion (MPC) [verbeke2012new], which estimates profit based on classification outcomes. This was extended by Verbraken et al. through the Expected Maximum Profit (EMP) metric [verbraken2012novel], which incorporates uncertainty in retention success and introduces the profit-maximising churn fraction, the optimal subset of customers to target. These metrics marked an important shift from accuracy-focused to business-aligned evaluation. Other approaches focus on value-aware targeting and campaign evaluation under operational constraints [rahman2025profit]. This approach propose fairness-aware preprocessing (e.g., reweighing and resampling) to improve representation and targeting high-CLV customers, and introduce expected profit-based evaluation across target fractions (the Area Under the Expected Profit Curve) to reflect budget-limited retention strategies. However, these value-aware approaches typically require modifying the training data or the learning objective (e.g., via reweighing/resampling), and their conclusions remain sensitive to how CLV segments and profit parameters are specified. Table 1 summarises the key characteristics of traditional and profit-aware evaluation metrics, highlighting the business relevance and flexibility offered by e-Profits compared to widely used alternatives.
@@
Nevertheless, limitations persist. EMP assumes constant values for customer lifetime value (CLV), contact cost, and offer acceptance probability [oskarsdottir2018profit]. It models retention success using an unimodal Beta distribution with fixed hyperparameters, which may not capture the heterogeneity of real-world customers. Additionally, while some studies have explored embedding profit functions directly into model optimisation via profit-aware decision trees or evolutionary strategies [stripling2018profit, hoppner2020profit]. For instance, ProfLogit replaces likelihood-based fitting with an EMP-based fitness function and uses evolutionary optimisation to search for model parameters that maximise expected campaign profit [stripling2018profit]. These approaches often introduce complexity and reduce interpretability, limiting their uptake in business environments.
@@
Some recent works have considered time-dependent churn risk using survival analysis models [perianez2016churn, xu2022lads]. These approaches explicitly demonstrate the value of incorporating tenure and retention time, but they seldom translate into actionable profit-based evaluation metrics.
@@
Metric Type Business-Aware Fixed Assumptions Individual-Level Optimisation Target Interpretability
AUC Ranking No No Churn likelihood High (threshold-free statistic)
F1-score Classification No No Class balance High (standard classification metric)
EMP Profit-based Partial Yes No Profit-max subset Medium (requires Beta and cost inputs)
AUEPC Profit-based Yes Partial Ranking-level only Area under profit curve Medium (campaign-level profit curve; customer heterogeneity enters via ordering)
e-Profits Profit-based Yes No Yes Total and segment profit High (transparent, per-customer profit decomposition)
While the academic literature contains several profit-aware metrics, their adoption in operational churn platforms remains rare. Surveys such as García et al. [garcia2017intelligent] and our own review [manzoor2024review] highlight that many practitioners still default to traditional metrics like AUC or lift, due to limitations in available tools and ease of understanding. This disconnect reinforces the need for practical and business-aligned evaluation metrics that can bridge the research-practice divide.
Marketing research has long emphasized that churn behavior is heterogeneous across customers and that ignoring such heterogeneity can lead to poor managerial decisions [fader2010customer]. In addition, churn risk is not equivalent to intervention responsiveness. Ascarza [ascarza2018retention] highlights that targeting customers solely because they are high risk can be ineffective if those customers are unlikely to change their behaviour in response to retention actions. This distinction is important for our setting. We therefore position e-Profits as a decision-oriented evaluation and model-selection criterion for a targeting policy: it maps a model’s churn scores to intervention decisions and aggregates the resulting costs and retained value under explicit business assumptions. Integrating heterogeneous responsiveness to interventions (e.g., uplift or treatment-effect modelling) is a complementary extension.
Despite these advances, existing profit-aware approaches often lack a simple and granular way to evaluate churn classifiers under customer-level heterogeneity in value, costs, and retention dynamics. Most rely on fixed cost parameters or make strong assumptions about offer acceptance rates. By contrast, our proposed e-Profits is aimed at leveraging individual-level cost, revenue, and survival-based retention to provide a metric that is granular, understandable, and practically deployable, without retraining the model. e-Profits responds to these limitations by offering a modular evaluation metric that aligns directly with business impact, empowering practitioners to make financially sound model selection decisions. In addition, we rigorously test whether resulting model re-ranking is statistically significant by reporting confidence intervals, non-parametric paired tests, and rank correlation analysis on held-out test data.
@@ 3.1 Overview and Motivation
As established in the previous section, conventional evaluation metrics fail to capture the financial implications of churn misclassification, and existing profit-based metrics often rely on strong population-level assumptions. In response, we introduce e-Profits, a modular evaluation metric that supports granular, business-aligned model comparison. The e-Profits metric evaluates model performance by aggregating per-customer profit outcomes, combining customer-specific lifetime value, tenure-conditioned one-period retention probabilities derived via Kaplan-Meier survival analysis, and individualised intervention costs including contact and retention offers. This design enables profit-based comparison of churn models without altering training pipelines, ensuring both understandability and practical deployment feasibility.
@@
Figure 1: Conceptual workflow of the e-Profits evaluation metric. This flow illustrates how model predictions interact with customer-level retention, CLV, and cost information to compute aggregate business impact. (To calculate CLV, we use ARR and TRR as one-period retention probabilities derived from Kaplan-Meier.)
@@
We define e-Profits as the total estimated profit that would be realised if a retention campaign was executed based on the model’s churn predictions. Consider a dataset of $N$ customers indexed by $i=1,\dots,N$. For each customer $i$ we denote by $y_{i}\in\{0,1\}$ the true churn label (1 indicates churn in the observation window), and by $\hat{y}_{i}\in\{0,1\}$ the binary churn prediction produced by a classifier at a fixed decision horizon. The term $\Pi_{i}$ represents the incremental profit or loss associated with customer $i$ under a retention policy that targets customers with $\hat{y}_{i}=1$. The proposed overall e-Profits is computed as the sum of individual customer-level profit contributions:
$\text{\emph{e-Profits}}=\sum_{i=1}^{N}\Pi_{i}.$ (1)
Eq. 1 states that e-Profits is an aggregate measure obtained by adding up $\Pi_{i}$ which is the estimated profit or loss associated with customer $i$. We use a fixed threshold of 0.5 to obtain binary predictions $\hat{y}_{i}$ from predicted probabilities.
3.2.1 Customer-Level Profit Function
For each customer $i$, profit is determined by the interaction between the true churn outcome $y_{i}\in\{0,1\}$, the model prediction $\hat{y}_{i}\in\{0,1\}$, the customer lifetime value $\text{CLV}_{i}$, and the cost of intervention. The per-customer profit is formally defined as:
@@ 3.2.1 Customer-Level Profit Function
This formulation reflects practical churn intervention logic: only customers predicted to churn receive retention actions. Each branch corresponds to one of the four possible combinations of prediction and true label for customer $i$. $\text{CLV}_{i}$ is the estimated customer lifetime value, and $C_{\text{offer},i}$ and $C_{\text{contact},i}$ are the intervention costs for customer $i$. When the model predicts churn ($\hat{y}_{i}=1$) and the customer is in fact a churner ($y_{i}=1$), we assume that the intervention prevents churn and the firm retains the estimated customer lifetime value $\text{CLV}_{i}$ but pays the offer and contact costs. This gives $\Pi_{i}=\text{CLV}_{i}-C_{\text{offer},i}-C_{\text{contact},i}$ for true positives. When the model predicts churn but the customer would not churn ($y_{i}=0$), the intervention does not increase revenue, yet the firm still incurs the offer and contact costs. The incremental profit is therefore negative and equal to $C_{\text{offer},i}-C_{\text{contact},i}$ for false positives. For false negatives and true negatives ($\hat{y}_{i}=0$), no action is taken and no retention cost is incurred. Since e-Profits evaluates the incremental profit of the decisions made under the targeting policy, we assign $\Pi_{i}=0$ in these cases. This makes eq. 2 a transparent mapping from classification outcomes to profit contributions. We emphasise that e-Profits evaluates the profitability of a targeting policy under explicit cost and retention assumptions; it does not estimate the causal treatment effect of an intervention.
3.2.2 Customer Lifetime Value (CLV)
Customer lifetime value for customer $i$ is computed by combining their revenue, a fixed profit margin and estimated retention probability:
$\text{CLV}_{i}=\frac{R_{i}\times M}{1-\min(r_{i},0.995)}$ (3)
Where $R_{i}$ denotes the monthly revenue of customer $i$, $M$ is a fixed profit margin, and $r_{i}$ is the estimated one-period retention probability (probability of remaining active for one additional billing period). The denominator $(1-r_{i})$ corresponds to the churn probability in a geometric-like revenue model, yielding the familiar perpetuity form commonly used in marketing and finance. Retention probabilities $r_{i}$ are obtained from a Kaplan-Meier survival estimator as a function of customer tenure, prioritising simplicity, low-overhead and transparency. To avoid unrealistically large CLV values when estimated retention probabilities approach one, we cap $r_{i}$ at 0.995 as a numerical safeguard; this cap does not represent a behavioural assumption. We emphasise that e-Profits evaluates the profitability of a targeting policy under explicit cost and retention assumptions and does not estimate causal treatment effects.
@@ 3.2.2 Customer Lifetime Value (CLV)
3.2.3 Intervention Cost Model
Retention Offer Cost
$C_{\text{offer},i}=\text{CPO}\cdot\text{CLV}_{i}$ (4)
where CPO denotes Cost-per-offer ratio (e.g., 0.1).
Customer Contact Cost
$C_{\text{contact},i}=\max\left(c_{0},\;c_{1}\cdot C_{\text{offer},i}\right)$ (5)
where $c_{0}$ represents a minimum fixed contact cost and $c_{1}$ scales variable contact cost proportionally to the offer cost. This parametrisation reflects the intuition that high value customers tend to receive more expensive offers and may be contacted through more resource intensive channels, while keeping the model simple and easy to calibrate. This structure reflects realistic business conditions where retention cost increases with customer value, while respecting minimum outreach overhead. Complete workflow of e-Profits is illustrated in figure 1.
3.2.4 Retention Probability Estimation via Survival Analysis
To estimate the probability that a customer remains subscribed beyond a given tenure, we employ Kaplan-Meier survival analysis [perianez2016churn]. This method is non-parametric and does not assume a specific distribution of churn times. In our setting, retention varies at the customer level only through each customer’s observed tenure using a single population-level survival curve. We therefore describe these quantities as tenure-based retention probabilities. The term tenure-conditioned retention emphasise that retention varies across customers only through current tenure $T_{i}$ via a single population Kaplan-Meier curve, rather than through a covariate-based individual survival model. We use the terms tenure-based retention and tenure-conditioned retention interchangeably to refer to retention probabilities derived from conditioning on current tenure via the Kaplan-Meier curve.
Train/test separation and censoring We fit the Kaplan-Meier survival curve $S(\cdot)$ using only the training split to avoid any information leakage into model evaluation. Churn events are treated as failures, and customers who have not churned by the end of observation are treated as right-censored at their observed tenure, under the standard non-informative censoring assumption. The time origin is subscription start (onboarding), tenure is measured in months. The held-out test split is never used to fit $S(\cdot)$.